OpenStudy (anonymous):

u + 5/4u + 3/5 = 3/5 (u - 1/8)

OpenStudy (anonymous):

9/5u+3/5=3/5u-3/40 6/5u=-3/40-24/40 6/5u=-27/40 6u=-135/40 u=-9/16

OpenStudy (anonymous):

just to clarify, is your question u + 5/(4u) + 3/5 = 3/(5 (u - 1/8)) or u + (5/4)u + 3/5 = (3/5) (u - 1/8)

OpenStudy (anonymous):

no the multiply is on the right side (u - 1/8) the first part is u + 5/4 a fraction + 35 a fraction + 3/5 a fraction times (u - 1/8 a fraction)